Find the​ z-scores for which 75% of the​ distribution's area lies between -z and z.
Answer #1

solution

middle 75% of score is

P(-z < Z < z) = 0.75

P(Z < z) - P(Z < -z) = 0.75

2 P(Z < z) - 1 = 0.75

2 P(Z < z) = 1 + 0.75 = 1.75

P(Z < z) = 1.75 / 2 = 0.875

P(Z <1.15 ) = 0.875

z  ±1.15